### v2

# The Composition and Chemistry of Titan’s Atmosphere 111Accepted for publication in ACS Earth and Space Chemistry

Abstract
In this article I summarize the current state of knowledge about the composition of Titan’s atmosphere, and our current understanding of the suggested chemistry that leads to that observed composition. I begin with our present knowledge of the atmospheric composition, garnered from a variety of measurements including Cassini-Huygens, the Atacama Large Millimeter/submillimeter Array (ALMA), and other ground and space-based telescopes. This review focuses on the typical vertical profiles of gases at low latitudes, rather than global and temporal variations. The main body of the paper presents a chemical description of how complex molecules are believed to arise from simpler species, considering all known ‘stable’ molecules – those that have been uniquely identified in the neutral atmosphere. The last section of the paper is devoted to the gaps in our present knowledge of Titan’s chemical composition and how further work may fill those gaps.
keywords:
American Chemical Society, LATEX+1 301 286-6757
NASA GSFC]
Planetary Systems Laboratory, NASA Goddard Space Flight Center
8800 Greenbelt Road, Greenbelt, MD 20771
[…]
1 Introduction
Titan, Saturn’s largest moon, was first observed in 1655 by the Dutch astronomer Christiaan Huygens. This important discovery that Saturn, in addition to Jupiter, had its own satellite helped to consolidate the Copernican worldview: that the Earth was no longer to be considered the center of the Solar System, but rather one of several planets orbiting the Sun, and possessing a natural satellite of its own. In his excitement at this new finding, Huygens would have little suspected what would transpire 350 years later: that a machine devised and launched into the heavens by humanity and bearing his name would traverse an unimaginable void, and then land softly on his new world, finding it stranger and more alien than even the machine’s designers had anticipated1.
Over the 13 years from 2004 to 2017, the Cassini-Huygens mission2, 3 was able to significantly reveal Titan, both to our eyes and to our minds. Between the successful landing of the ESA-built Huygens probe carrying six scientific suites in January 2005, and the 127 flybys of the NASA-built Cassini spacecraft with its own twelve science instruments, our knowledge of Titan now is vastly greater than before the mission arrived. So, in our privileged position of hindsight, what do we now know about Titan, almost four centuries since its discovery?
First and foremost, it is a moon with a dense atmosphere (Fig. 1), the only such body known in our Solar System. Also that this atmosphere, composed primarily of molecular nitrogen and methane, is a largely anoxic environment, with little oxygen to cause the termination of complex organic reactions. The result is a chemical wonderland, with a breathtaking array of complex organic molecules, of which we presently have only the most rudimentary understanding. Fig. 2 shows a schematic overview of the presumed chemistry that occurs in Titan’s upper atmosphere, where the ‘raw ingredients’ of its photochemical reactions, N and CH, are broken apart and recombined into successively larger molecules, and finally haze particles.
Titan is a world that is today both tantalizingly more known, and more unknown than ever before. Our direct investigations of its atmosphere by the Huygens lander and Cassini spacecraft have both increased our understanding of Titan enormously and also multiplied our questions. Significant outstanding questions include:
Why is Titan the only moon in the Solar System with a significant atmosphere, and how did it come to be in its present state?
Is the atmosphere today in a steady state, or is it growing or shrinking or changing in some way?
To what degree has the atmosphere interacted with and shaped the surface and subsurface?
What degree of chemical complexity is reached in Titan’s atmosphere, and are precursor biomolecules among the products?
Are there other moons with atmospheres similar to Titan elsewhere in the galaxy?
Achieving a better understanding of Titan’s atmosphere and its chemistry is important both for the sake of Titan science itself, and because of its potential to inform us about other environments. This includes the present-day Earth, since Titan and Earth are the only objects in the Solar System today to have a hydrological cycle of evaporation, condensation and precipitation, and associated rivers, lakes and seas 4, 5, 6, 7. Like the Earth, Titan also experiences seasons, due to orbiting close to Saturn’s equatorial plane, which is tilted 27 to the ecliptic. Titan therefore experiences summer and winter in each hemisphere, seasons that last 7.4 longer than on Earth, with transitional equinox periods of equal daylight at all latitudes.
We also note the relevance to the early Earth, which likely had a much more chemically reducing atmosphere in its distant past 8, 9, 10, 11, 12, 13, before the Great Oxidation Event 14, 15, 16. Finally we can surmise the likely relevance to exoplanets, which vastly outnumber the planets in our own Solar System, and more likely than not include Titan-like bodies somewhere in our galaxy 17, 18, 19.
In this review paper, I attempt to lay out a simple picture of the known characteristics of Titan’s atmosphere, with a focus on the composition and chemistry of the dense lower atmosphere. By necessity, this review will not cover, except in passing, many related areas: the origin of the atmosphere and possible replenishment mechanisms by internal or external sources; isotopic composition and time evolution of isotopic ratios; winds and dynamics; condensates and meteorology; and the chemical composition of large particulates (haze particles). All of these topics have been covered extensively in review articles and chapters elsewhere7, 20, 21, and in two books written about the results of the Cassini-Huygens mission 22, 23.
In presenting a simple overview and summary focusing on the chemistry and composition of the neutral atmosphere, it is hoped that I will do sufficient justice to this one area to make this article a useful primer for undergraduate or graduate students, or others new to the field, to quickly gain a basic understanding of Titan’s bulk atmospheric composition, and why it is that way - at least at the present era.
The paper is organized as follows: I first review basic knowledge about Titan’s atmospheric temperature structure and gas composition. The main section of the paper contains an exposition on the chemistry of the 24 known molecules in the neutral, lower atmosphere. This is followed by a detailed discussion of future research directions in Titan atmospheric composition studies, followed by a Summary and Conclusions.
2 Atmospheric Composition and Structure
2.1 Atmospheric Composition
Titan’s atmosphere is largely composed of two gases: N and CH. The vertical profile of methane comes from measurements by instruments on the Cassini-Huygens space mission, primarily the Huygens GCMS (Gas Chromatograph and Mass Spectrometer) 24 from 0–146 km, the Cassini Visual and Infrared Mapping Spectrometer (VIMS)25 (50–850 km), the Cassini UVIS (Ultraviolet Imaging Spectrometer) 26 (400–1650 km), and the Cassini INMS (Ion and Neutral Mass Spectrometer) 27 (900–1500 km). Their results have been reported in publications from the mission 28, 29, 30, 31.
Aside from noble gases (Ar, Ar and Ne) 32, 31, 22 molecular species other than N and CH have been definitively detected in Titan’s atmosphere at the time of writing (see Fig. 3): ten hydrocarbons (CH, CH, CH, c-CH, CHCCH, CHCCH, CH, CH, CH, c-CH), eight cyanides222Also referred to as ‘nitriles’ when -CN occurs in an organic molecule. (HCN, HNC, HCN, CN, CHCN, CHCN, CHCN, CHCN), three oxygen-bearing species (CO, CO, HO) plus H. These gases were originally detected by a variety of astronomical and remote sensing techniques from the ground and space.
All of the major types of hydrocarbons have been detected (alkanes, alkenes, alkynes, a carbene, an aromatic ring). However, major chemical families of nitrogen-bearing molecules (including amines, imines, azines, and N-heterocyclic rings) and oxygen-bearing molecules (such as aldehydes, ketones, alcohols, ethers) are possible ingredients of the atmosphere but remain undetected - a subject we will return to in a later section.
Oxygen has yet to be detected on Titan in an organic molecule such as methanol (CHOH) or formaldehyde (HCO), being found so far only in the simple inorganic molecules CO, CO, and HO. This limits the presently confirmed scope of astrobiological molecules (i.e. those with the elements CHON in a variety of functional groups) - at least in the atmosphere. At the surface and in the subsurface - where hydrocarbons are thought to be readily hydrolyzed as seen in laboratory experiments 33, 34, 35 - the astrobiological potential may be much greater 36, 37, 38.
The reaction pathways that lead between these molecules have been compiled into computational models of the atmospheric chemistry, which have largely been successful at replicating the observed gas abundances. Models pre-dating the Cassini-Huygens mission 39, 40, 41, 42, 43, 44, 45, 46, 47 primarily focused on replicating the observed neutral gas abundances as measured by Voyager 48 and the Infrared Space Observatory (ISO) 49. However, some models were also developed for the ionosphere 50, 51, 52, 53, 54. During the Cassini-Huygens mission and since, new information collected by the spacecraft, especially from direct sampling of the ionosphere,28, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78 has prompted many new and revised models of Titan’s atmosphere 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103.
At the opposite end of the size scale, molecular growth by covalent bonding and agglomeration results in macro-molecular haze particles 104, 105, 106, 107, 108, 11, 61, 109, 110, 111, 112, composed of thousands to millions of individual atoms 113. As these particles reach a size of 1 m, they begin to sediment (or form the nuclei for condensate growth) and are removed from the atmosphere 114, 115, 116, 117, apparently forming vast dune fields on the surface 118, 119, 120.
2.2 Atmospheric Temperature Structure
Titan’s atmospheric temperature structure (Fig. 4) is a result of the competing heating and cooling processes that take precedence at different altitudes. In the dense lower atmosphere, convection driven by surface heating leads to a vertically decreasing temperature profile, as warm air rises and adiabatically cools. A temperature minimum is reached at 45 km, the tropopause, as confirmed by direct measurements121 from the Huygens Atmospheric Structure Experiment 122 (HASI) and occultations by the Cassini Radio Science Subsystem123 (RSS) 124, 125, 126.
Above the tropopause, heating by absorbed solar energy, primarily by atmospheric haze particles, causes temperatures to rise again in the stratosphere 129. Stratospheric temperatures have been measured by Cassini HASI 121 and by the Cassini Composite Infrared Spectrometer 130 (CIRS) 131, 132, ALMA 133, 134, 128, as well as through radio occultations for the lower stratosphere 124, 125, 126. At around 250–400 km, a temperature maximum, the stratopause, is reached at 180 K 135. The exact altitude (pressure) and temperature of the stratopause varies with both latitude and season 136, 137, 127 over the course of Titan’s long year (29.46 Earth years), being higher and warmer (by 20 K) over the winter pole. This somewhat counter-intuitive result can be understood as resulting from adiabatic compression of air in the descending branch of the global stratospheric Hadley cell.
Temperatures fall throughout the next layer, the mesosphere, as haze becomes thin, and radiative cooling by gases such as HCN and CH becomes increasingly important 129, 135. Titan’s mesopause is reached at 600 km 128, above which altitude temperatures rise again. This is primarily due to methane UV absorption 138, blocking of outgoing IR radiation by CH, and far-infrared HCN rotational lines 129. This is the thermosphere, a region where gas collisions are rare, and molecules must wait to spontaneously emit a photon to lose energy.
[…]
Above the four layers of the bound atmosphere is the exosphere, beginning at the exobase (1500 km 148), a region where gases can freely escape to space. These five regions mirror the temperature structure of the Earth’s atmosphere, but with a substantially larger scale height (approximately 5) due to the lower surface gravity. Overlapping the upper thermosphere is the ionosphere ( km), defined as the region where "significant numbers of free thermal (1 eV) electrons and ions are present." 149
It is important to note that the vertical profiles of temperature, minor gas abundances and haze density all vary with both latitude and season. Titan - like the Earth and other planets with atmospheres - exhibits one or more convection cells in the middle atmosphere. Near to the equinoxes, air rises at mid-latitudes and flows to both poles, where it descends and then returns equator-ward 150, 151, 152, 153, 154. However, close to the solstices the circulation more closely resembles a single cell with flow from the summer to winter hemisphere. These cells act to redistribute thermal energy, trace gases and hazes in both altitude and latitude. This topic has been the subject of extensive measurements and modeling in the literature (e.g. 127, 155, 156, 157, 158, 159, 160, 161, and references therein). In this review I will not further discuss latitudinal or longitudinal variations in atmospheric structure and focus only on the vertical chemistry variations typical of mean conditions at low latitudes.
2.3 Gas Vertical Profiles
Fig. 5 shows typical vertical profiles of the 24 known molecular species at low latitudes, compiled from a combination of ground and space-based measurements, and some photochemical model profiles constrained by observations. These include two pairs of structural isomers: HCN and HNC, CHCCH and CHCCH.
Some important trends can be noted. Methane is an unusual outlier, with a greater abundance in the troposphere ( km) than above. This is due to the ‘cold trap’ effect, where it reaches saturation as the tropospheric temperature drops with altitude, and therefore its mixing ratio is reduced as it forms clouds at 15-30 km. The gases N and CO are well-mixed, having approximately uniform profiles throughout the atmosphere due to long photochemical lifetimes. Two other gases: H and CH also do not condense at the ‘cold trap’ - the coldest part of the atmosphere around the tropopause at 45 km (70 K). The profile of H is shown here as constant at the 0.1% value typical of the lower stratosphere, since measurements of its vertical profile remains uncertain.
All other gas species show profiles that typically decrease downwards from the upper atmosphere, due to having a source due to photochemistry at high altitudes, and then becoming diluted as they are mixed downwards into the denser part of the atmosphere. In many cases the actual measured profiles are still rudimentary, constrained by only a few data points and with even less knowledge of meridional and temporal variations. Some gases may exhibit increases again towards the stratosphere, due to either secondary production peaks (e.g. due to cosmic ray deposition) or due to redistribution by atmospheric circulation.
3 Atmospheric Photochemical Processes
In this section I present a brief overview of the types of reactions that occur in Titan’s atmosphere, as a prelude to the discussion of the chemistry of individual molecules in the following section (see also Table 5 of Vuitton et al. 103 and description therein). Different chemical processes become important at different levels of the atmosphere, due to the altitude variation of temperature, density and penetration depths of charged particles and photons that affect the reactions. For example, Saturn magnetospheric electrons are stopped high up in the ionosphere, while solar photons penetrate to varying depths depending on wavelength.103 High energy cosmic rays are the deepest-penetrating rays in the atmosphere, peaking in energy deposition at around 100-150 km altitude.103 The various processes, depicted in Fig. 6, are now examined in more detail.
3.1 Photodissociation and electron impact dissociation
Dissociation (break-up, or fragmentation) of a molecule occurs when a sudden influx of energy breaks molecular bonds. In Titan’s atmosphere this happens readily in the upper atmosphere (700 km), due to both energetic UV solar photons (h) and by the impact of fast-moving electrons () trapped in Saturn’s magnetic field 172, 173 (Fig. 2).
Dissociation often leads to neutral (uncharged) molecular fragments, which may either be in the ground state or excited states, e.g.:174, 175, 176
where is ground state methylene and is an excited state. Note that other possible fragmentation products are possible, the examples given above are only one possibility in each case. Also note that for simplicity, electronic states of molecules are usually omitted in this paper, unless required to distinguish between two otherwise identical reagents that have significantly different properties.
The molecular fragments are often radicals – i.e. highly reactive atomic or molecular species that have unpaired electrons, such as H, CH, and N. These radicals are quick to react with other radicals, or with neutral species.
3.2 Ionization
Instead of breaking up (dissociation), a molecule may instead become ionized (positively or negatively charged), typically by losing an electron to become a cation, e.g.: 181, 182, 29
3.3 Ion reactions
Dissociative recombination is the process whereby a positive ion reunites with an electron, and in the process breaks apart. An example in Titan’s atmosphere is: 188, 189
Radiative association is a reaction whereby two species combine, shedding excess energy via a photon. These reactions typically occur only in rarefied environments ( mbar190 where a metastable intermediary complex has time to form and then stabilize by emission of a photon. Despite being impossible to observe in laboratory conditions due to the long lifetimes of the intermediate states191, 192, such reactions are thought to occur in interstellar clouds193, 194, 195, 196, as well as in Titan’s upper atmosphere. Radiative ion-neutral association reactions on Titan may include: 197, 103
[…]
3.4 Radical reactions
Radicals (molecules with an unpaired electron) react with other radical and non-radical species in multiple ways. Radicals may react with each other in association reactions 47:
Radicals may attack neutral molecules, for example in the case of hydrogen abstraction, which is a major loss process for methane 41, 199, 200, 103:
Other examples include substitutions of terminal atoms or groups, typically at carbon-carbon double or triple bonds 201, 202, 203:
Three-body association reactions occur when two reactants meet to form a metastable, intermediate complex that is then stabilized by collision with a third, non-reacting body that carries away energy, allowing the metastable complex to stabilize. The two steps are:
where the asterisk is used to denote an excited state. In this paper I typically simplify such reactions to a single step:
Such reactions are of critical importance to the formation of many hydrocarbons, especially alkanes, e.g.207, 208, 197:
Note that three-body reactions are limited to Titan’s dense, lower atmosphere where there is a sufficiently high collision rate to allow the collisional stabilization to occur.94
Radiative association reactions may also occur between radicals. Vuitton et al.94 studied the effect of radiative associations on Titan’s chemistry and proposed that reactions such as:
may occur in Titan’s atmosphere.
3.5 Other reactions
Molecules may also re-organize their structure to become more stable, such as in collisional isomerization. Two important known instances of this on Titan are the conversion between propadiene (allene) and propyne by H atoms,41, 99 with a barrier of 65.1 kcal/mol:209
Polymerization is the process where multiple similar, unsaturated hydrocarbons join together form linear chains:
4 Chemistry of the Neutral Atmosphere
[…]
Note on chemical names: the topic of chemical nomenclature remains eternally problematic. For example, the relatively simple compound CHCN has been commonly referred to as ‘vinyl cyanide’ in most 20th century astronomical literature, though preference has shifted more recently towards the simpler, single-word name ‘acrylonitrile’ (also with ‘methyl cyanide’ to ‘acetonitrile’). In fact multiple valid names for CHCN exist, including ‘2-propenenitrile’, ‘cyanoethene/ cyanoethylene’ and ‘propenenitrile’, however it must be noted that the preferred official (IUPAC) name is actually the rather cumbersome ‘prop-2-enenitrile’.
IUPAC molecular nomenclature certainly has its place. However, for the purposes of discourse in planetary atmospheric chemistry, dominated by small molecules, the formulaic names for molecules can be not only inconvenient, but an actual obstacle to reading and digesting information. Therefore, in this work I have followed a naming convention based on a combination of modernity, modified in places for greater simplicity, i.e.: (a) single word names are generally preferred over multiple word names (e.g. ‘propyne’ over the older standard ‘methyl acetylene’); (b) avoiding numbers in names of small molecules except where necessary, and (c) accepting that some names are close enough to be interchangeable (e.g. there is little confusion engendered by using either ‘ethylene’ and ‘ethene’; ‘propylene’ and ‘propene’ etc).
While I run the risk of offending practicing chemists, I hope that the terminology is consistent enough for readers to understand what molecule is being referred to, and convenient enough for simplified writing in typical planetary science usage.
4.1 Hydrocarbons and Hydrogen
Hydrocarbons are molecules formed from atoms of only hydrogen and carbon. Due to the four-fold valency of carbon, many bonding configurations are possible. The most common families of aliphatic (acyclic) hydrocarbons include (i) the alkanes - where carbon is saturated having four single-bonds; and the two unsaturated types - (ii) alkenes, featuring carbon-carbon double bonds, and (iii) alkynes with carbon-carbon triple bonds. Once there are four or more carbon atoms, mixed types become possible (see Fig. 8).
Cyclic hydrocarbons occur where there are one or more closed rings of carbon atoms (at least three are needed to make a ring, or cycle). Cyclic molecules that have a pi bond of delocalized electrons are known as ‘aromatic’ (small aromatics are typically volatile at room temperature), although not all rings are aromatic. For example the unsaturated six-carbon ring benzene (c-CH) is a well-known aromatic, whereas its saturated cousin cyclohexane (c-CH) is a non-aromatic cycle. Carbenes are reactive molecules where carbon has two unbonded but paired electrons, such as methylene (CH).
Hydrocarbon molecules are created through the breakup and recombination of fragments of methane. A prime example is the formation of ethane from two methyl radicals:221, 222
Note that, in the process, significant amounts of molecular hydrogen will form from the photo-dissociated hydrogen atoms 223:
leading to the trace amounts of hydrogen (0.1%) found in Titan’s lower atmosphere. Note that this termolecular reaction is an important process leading to the creation of H in the Interstellar Medium (ISM).
Significant loss of hydrogen to space is thought to occur,40, 224, 148, 225, 226, 227, 228 preventing methane from being recycled, as occurs on the giant planets, leading to gradual depletion of methane in Titan’s atmosphere in the absence of outgassing or other replenishment.41, 229, 230, 231, 232
A network diagram showing the principal neutral pathways for hydrocarbon molecule formation is shown in Fig. 9. Hereafter follows a high-level description of the key chemistry for each of the hydrocarbons and hydrogen leading to their relative abundances in the neutral atmosphere.
4.1.1 Hydrogen
Detection: Molecular hydrogen (H) was was tentatively detected by Trafton233 using the 107 inch telescope at the McDonald Observatory, via absorption in the S(0) and S(1) quadrupole lines of the 3-0 band at 0.82 m. Hydrogen was clearly detected in the far-infrared by the Voyager 1 IRIS spectrometer 234 and confirmed by Cassini CIRS 235. The H Volume Mixing Ratio (VMR) was measured directly by the Cassini INMS instrument in the ionosphere as 0.4% 28 and in the lower stratosphere and troposphere by Huygens GCMS at 0.1% 31. The fact that hydrogen, presumed to be produced in the upper atmosphere by photolysis of methane 40, 41, was measured to have a decreasing abundance downwards, has proved difficult to replicate in models. Models have required a sink for H at the surface 225, 135, 236, which has even been suggested as possibly biological in origin 237. Important reactions for hydrogen are shown in Fig. 10 238.
Production: Molecular hydrogen can be produced in ion-phase reactions such as239:
[…]
4.1.2 Methane
Profile: Methane (CH) was the first molecule to be positively identified in Titan’s atmosphere, via visible and near-IR absorptions seen by Gerard Kuiper 246, using the 82 inch reflector at McDonald Observatory. We now know that methane is the basic ingredient enabling all of Titan’s complex organic (i.e. carbon) chemistry, allowing reactions to proceed up to the creation of haze particles: some key reactions for methane are shown in Fig. 11247, 248.
Nevertheless, methane may be a gradually depleting resource since it is not permanently recycled, unless replenished by an as yet unidentified mechanism 41, 249, 232, 231, 250. Speculative mechanisms include: crustal destabilization leading to outgassing from methane clathrates,251 outgassing from cryovolcanism,252, 253, 254 and displacement from near-surface clathrate materials by condensed ethane255 - yet observational evidence for these processes remains inconclusive at best.
Methane’s vertical profile can be divided into three zones: (i) a tropospheric zone where the fractional abundance gradually decreases from 5.5% at the surface to a minimum at the tropopause of around 1.4%, due to reaching saturation at decreasing volume mixing ratios (VMRs) as the temperature decreases towards the tropopause (‘cold trap’); (ii) a relatively constant amount of 1.4% in the stratosphere, mesosphere and thermosphere; (iii) a gradually increasing mixing fraction above an altitude of 800–850 km (the methane homopause) in the increasingly collisionless regime, due to the differing scale heights of different molecules 135. It is presently uncertain whether methane has any variation with latitude on Titan, although a variation from 1.0% to 1.5% in the lower stratosphere has been reported 256 based on infrared measurements by Cassini CIRS.256
Loss: The photolysis of methane in the upper atmosphere leads to the formation of radicals including methyl (CH), methylidene (CH), and the carbene methylene ( or )103 which undergo a chain of reactions to form all the hydrocarbons found in Titan’s atmosphere. Note that as much as 75% of methane photolysis above 700 km is due to the solar Lyman- line at 121.6 nm.257
The fate of most methane is ultimately to form ethane via the addition of two methyl radicals (R 31) or via the creation of ethyl:
leading to permanent loss of methane. Methyl radicals are produced either directly by primary photolysis, or by reaction of methane with radicals:205, 258
Note that the reaction of the ethynyl and vinyl radicals (amongst others) with methane is a catalytic destruction process, since the acetylene and ethylene generated are easily photolyzed back to the radical form where they can continue to destroy methane molecules. This process may be repeated hundreds of times before the radical catalysts are themselves lost to form higher hydrocarbons. This is the main source of methane depletion in Titan’s atmosphere.103
Methane fragments participate in ion chemistry, first being ionized by charge transfer:242
and then building up ions, e.g.:239
Future directions: Although the chemistry of methane is perhaps one of the best-understood for any molecule on Titan, the most pressing questions remain the nature of its origin, and possible replenishment.21
4.1.3 Acetylene
Acetylene was the third molecule to be identified in Titan’s atmosphere,259 following the detection of methane and ethane, when Gillett observed its 13 m band in mid-infrared spectroscopy with 2 m and 4 m telescopes at Kitt Peak Observatory. Important reactions for acetylene are shown in Fig. 12.260, 261, 257, 103
Production: Acetylene may be produced either directly or indirectly from photolysis products of methane. In a direct process262:
Loss: In the ionosphere, photolysis of acetylene, and electron transfer to N produces 242, which reacts with neutrals to build heavier ions239, e.g.:
which may (dissociatively) recombine with to form neutral species266, 267. Likewise, proton transfer to neutral acetylene leads to , which can also form species239:
Finally, neutral CH in the ionosphere may combine with other ions to build heavier species239, e.g.:
In the neutral atmosphere, acetylene is lost by reaction with methylene forming propargyl268, 269, 270:
Acetylene absorbs photons to longer wavelengths (230 nm79) than methane, so its photolysis continues into the stratosphere (see Fig. 12). This produces ethynyl (CH) and carbyne (C) which are potent means of methane depletion via hydrogen abstraction:223
The acetylene produced by this reaction is recycled back to ethynyl by photolysis, and thereby each acetylene/ethynyl may cause the loss of hundreds of methane molecules before being lost itself to another reaction pathway such as:271
Catalytic destruction of methane in this way is the principle means of methane depletion in Titan’s atmosphere 41, 79, 103. In the lower atmosphere ( km), the dominant loss process for acetylene is conversion to ethylene by a two-step reaction with atomic hydrogen:94, 103
Future directions: Acetylene has been detected on Titan’s surface 32, 272, and is likely to be present in the northern lakes and seas 273, 274. Future investigation by the mass spectrometer (DrAMS) instrument of Dragonfly275 will further refine the surface and near-surface abundance.
CH was one of the first molecules investigated to form a co-crystal 276, 277 at Titan surface temperatures, an organized co-condensate of two or more chemical species. The validity of multiple co-crystal types has since been established, but requires further laboratory work to determine the full parameter space of possible crystalline types.
4.1.4 Ethylene
Ethylene (CH) was discovered by infrared spectroscopy at the same time as acetylene 259. The vertical profile of ethylene exhibited a surprising trend to decrease in abundance upwards in the lower stratosphere early in the Cassini mission163, 278, although this faded at later seasons 137, 127.
Ethylene is a crucial, two-carbon neutral molecule that provides a stepping stone from methane to higher hydrocarbons (Fig. 13)279, 280. Ethylene is remarkable in being one of the few molecules (along with H and N) that does not condense at the tropopause, and therefore persists in significant quantities into the troposphere.
Production: Ethylene may be produced in the ionosphere by dissociative recombination of heavier ions with an electron281, 282, 266:
In the neutral atmosphere, ethylene is largely formed through reactions between methane and its derived radicals, or between radicals283, 257, 262, 79, 268, 269:
At lower altitudes, production through the intermediary is also important223:
Loss: In the ionosphere, the ethylenium ion - produced from ethylene photoionization - can be lost in various reactions with neutrals239:
and ethylene can be lost during the formation of heavier ions239:
Photolysis of ethylene leads to acetylene (R50 and R51). Insertion/addition reactions onto ethylene by CH, for example, can lead to higher hydrocarbons283, 204, 284:
Below 500 km, loss via H-addition becomes important94:
The CN radical can also substitute onto ethylene to create vinyl cyanide (also known as acrylonitrile, or propenitrile):98, 202, 203
Future directions: Ethylene is the simplest alkene, the family of hydrocarbons having a double CC bond. Photolysis or other breaking of the alkene CC bond leads to radicals which rapidly react, leading to formation of polymers. The role of polymer formation in Titan’s atmosphere is incompletely understood, but is likely to be an important process in the formation of haze particles.
4.1.5 Ethane
Ethane was the second molecule to be discovered in Titan’s atmosphere, via its strong band at 12 285 seen by Gillett with the 60-inch telescope on Mount Lemmon. Ethane forms one of the primary trace gases in Titan’s atmosphere with concentrations greater than 1 ppm in the stratosphere 48, 286, 287, and is the primary sink for methane loss 41, 47. Due to this observation, a global deep ethane ocean was originally predicted 288, 289 but later proved not to be the case 290, 291.
The liquid hydrocarbon bodies eventually detected on Titan’s surface292 have a measured ethane content that varies between different seas and is generally less than the methane fraction 293, 294, except for the southern lake Ontario Lacus 295, 296. Ethane may form co-crystals in Titan lakes with other organics, such as benzene 297. Ethane has also been implicated in displacing methane from clathrate hydrate, allowing for a partial resupply mechanism of methane to the atmosphere 255, which is otherwise continuously lost by chemistry 41, 298.
Production: Ethane is primarily produced by the addition of two methyl reactions (R31) but also reforms from the ethyl radical299:
Loss: Ethane can be lost through photolysis back to , or to stable molecules such as ethylene and acetylene with loss of hydrogen (see Fig. 14). Ethane can also be attacked by reactive radicals such as methylene, methylidene and ethynyl (from acetylene photolysis)223, 79, 283, 204:
Ethane may form heavier ions through reactions such as239:
Future directions: Large amounts of ethane are thought to condense in Titan’s lower stratosphere, and form a significant fraction of Titan’s lakes and seas 273, 303. Ethane was implicated in the formation of a vast north polar cloud seen during northern winter in 2005 by Cassini VIMS 304, although other interpretations have suggested that this cloud is condensed methane.305 While the basic chemistry of ethane is well-understood, an improved understanding of its condensation - especially co-condensation with other gases - will be crucial to a more accurate interpretation of Titan’s meteorology.
4.1.6 Cyclopropenylidene
Cyclopropenylidene (c-CH) is the first carbene (a molecule having two unbonded, self-paired valence electrons from a carbon atom) and the second cyclic molecule to be found in Titan’s atmosphere (after benzene). c-CH was detected using millimeter wavelength astronomy with ALMA 165, the third molecule whose first detection on Titan was achieved with with this telescope.
Production: In the upper atmosphere, production of a precursor, the cyclopropenyl cation () is thought to proceed by103:
which then recombines with to produce c-CH 306 and H.
Other possible pathways include photolysis of propargyl:
In the neutral atmosphere the dominant production pathways may include CH addition to acetyene 307, 308 (see Fig. 15)309, 310:
and below 600 km:311
c-CH can also result from collisional isomerization from its isomers propynylidene (t-CH) and propadienylidene (l-CH):311
which in turn may dissociatively recombine with an electron, returning to c-CH, or splinter into smaller acyclic fragments:266, 313
In the neutral atmosphere above 600 km, c-CH photodissociates to form c-CH, l-CH, and C. It is also lost by reaction with CH to form acylic species:311
Below 600 km, these pathways continue to be significant, but there is additional c-CH loss via:311
Future work: Chemical pathways leading to and from c-CH in Titan’s atmosphere remain to be explored, especially given the multiple possible structures for , including propynylidene ( ), propadienylidene (), cyclopropyne () and propenediylidene () 309, 310. Experimental and theoretical work on reaction pathways, rates and branching rates during photolysis will greatly help to clarify the production, loss and stability of cyclopropenylidene.
Several isomers of have been detected in space, including HCCC 314, which should prompt further astronomical observations to determine if these isomers also exist in Titan’s upper atmosphere. Intriguingly, the CHCCH form has been shown to dimerize to form para-benzyne 307, discussed in a later section.
4.1.7 Propyne
The propyne (CHCCH, methyl acetylene) isomer of () was first detected in the infrared by the Voyager Infrared Interferometer Spectrometer (IRIS)315 via long wavelength infrared emission bands at 328 and 633 cm316.
Production: Propyne and its symmetric isomer propadiene (CHCCH, allene) are produced by addition of CH into ethylene (Eq. 79, Fig. 16)317, 318, 319, but also through H-addition to propargyl320:
In the ionosphere, the CH ion is a precursor to CH, and produced via:219
Loss: In the ionosphere, propyne is lost through ion reactions such as239:
Propyne may also be lost by photolysis in the upper atmosphere324:
and to three-body reactions103:
Future work: Many reactions forming or depleting have uncertain branching ratios between CHCCH and its isomer CHCCH. Further work is needed to improve knowledge of these quantities. Collisional interconversion between the two isomers may be mediated by atomic hydrogen99, so accurate measurement of both isomers may be a way to provide constraint on the abundance of otherwise short-lived and difficult to measure H atom.
4.1.8 Propadiene
Propadiene is a less abundant and less thermodynamically stable isomer of CH, which is more abundant in Titan’s atmosphere in the form of propyne. Propadiene (CHCCH) was detected in Titan’s atmosphere using high-resolution ground-based spectroscopy at NASA’s Infrared Telescope Facility (IRTF) with the Texas Echelon Cross Echelle Spectrograph (TEXES) instrument via its band at 845 cm 325.
Production: Like propyne (CHCCH), propadiene is produced in the upper atmosphere by CH addition to ethylene (Eq. 79, Fig. 17)326, 318, 327, 328, and by H addition to CH:320
Lower in the atmosphere, where propene is more plentiful, it can be photodissociated to produce propadiene (see Fig. 18)329:
Ion formation pathways of CHCCH are less certain, but may follow similar channels as propyne, with branching ratios that are currently uncertain.
Loss: Propadiene is lost to direct photolysis318:
to ion reactions such as103:
Future directions: As with propyne, the branching ratios of reactions implicated in the formation of isomers remain uncertain, so further experimental and theoretical work is required. Accurate measurement of both propyne and propadiene is a possible means to indirectly inferring the abundance of atomic hydrogen99 in the lower atmosphere, in the absence of direct, in situ measurements.
4.1.9 Propene
Propene (CH, propylene) was first detected using data from the Cassini CIRS infrared spectrometer via its band emission near 11 m 330, 171.
Production: Ion reactions lead to creation of propene from , :239
Propene is predicted to be produced in the upper atmosphere by both H addition to (62%), and CH insertion into ethane (38%)91:
and also lower in the atmosphere by the termolecular reaction103:
Loss: Propene is lost both through photodissociation (Fig. 18331, 332, 333, 334, 321, 329) and ion reactions e.g.:239
In the lower atmosphere, propene is predicted to be lost through a termolecular reaction with H atom addition:103
Future directions: Propene, as an alkene, may also undergo polymerization to form polypropylene, a notable and widespread plastic used on Earth. Most likely polyynes in Titan’s atmosphere are not pure polymers of a single repeated monomer type (ethylene, propylene, etc) but an assorted mixture of many types, with the lighter, more abundant alkenes more heavily represented than larger, heavier units. Further research into polymerization of mixed monomers will yield insights into the formation of Titan’s haze.
4.1.10 Propane
Propane (CH) was detected contemporaneously with propyne (CHCCH) by Voyager’s IRIS instrument 316 via an infrared band at 748 cm and subsequently confirmed by ground-based observations335 and with Cassini CIRS 286, 163, 336.
Production: A significant pathway for the production of propane is by addition of to (Fig. 19)337, 338:
Loss: Propane is primarily lost in the upper atmosphere by photolysis to propene, but participates in other reactions as shown in Fig. 19. Propane also undergoes H-abstraction by ethynyl to recycle acetylene:340
but the fate of is largely to react with H to reform propane.103
Future directions: While the chemistry of propane remains relatively well known, its role in cloud formation and lake composition on Titan remains to be fully explored. Quantum mechanical analysis of propane’s 23 infrared active bands336 remains incomplete, preventing accurate modeling at high resolution for these bands. However in recent years the pseudo-linelist technique has proved useful for providing practical absorption coefficients across a wide bandwidth for calculation at medium resolution 341.
4.1.11 Diacetylene
The presence of diacetylene (CH, butadiyne) was inferred from infrared spectroscopy of Titan’s atmosphere with Voyager’s IRIS instrument 342 via emission bands at 220 and 628 cm. At present, it remains the only C hydrocarbon species confirmed in Titan’s atmosphere (although note that the nitrile CHCN, detected with ALMA168, also has four carbon atoms.)
Production: Diacetylene can be produced by the aforementioned reaction of the ethynyl radical with acetylene (R 63), or by stepwise addition to acetylene 283, 343, 103:
Loss: Diacetylene may undergo ionization to (Fig. 20)344, 345, and subsequent loss to processes such as:239
Insertion of ethynyl is a way to lengthen the polyyne chain from diacetylene to triacetylene:347
Future directions: To date, the triacetylene molecule () has remained elusive on Titan, despite its detection in space at high relative abundances compared to CH 348. Detection of triacetylene would help to clarify the importance of the radical, which contributes to the depletion of methane in photochemical models along with the smaller related radicals and , as well as the efficacy of polyyne formation in general.
4.1.12 Benzene
Benzene (c-CH, Fig. 21)349, 102, 350, 351, 352 was the second new species detection on Titan made by the Infrared Space Observatory (ISO) in 2003 49, via its strong hydrogen bending mode at 674 cm. Benzene was the first cyclic (closed ring) molecule to be detected on Titan, and remains the only confirmed aromatic molecule (molecules with a de-localized electron orbital). The detection of benzene is highly significant, since it provides a measurement of the basic six-membered ring from which larger, multi-ring molecules can be formed,353, 354 building towards macromolecular haze particles - see discussion later in this article.
Production: In the upper atmosphere (800-950 km), a significant pathway to creation of benzene is dissociative recombination (DR) of the phenylium ion. Phenylium is created through:92, 102
which then forms benzene through:355
A second source roughly equal in importance is thought to be dissociative recombination of the CH ion (benzylium or tropylium, Fig. 22):102
followed by:
Yet a third ion channel is the DR of CH with leading to benzene plus other hydrocarbon fragments.102
Radical chemistry also leads to to benzene, such as C addition to 1,3 butadiene proposed to occur in the ISM:356
An alternate pathway:
also leads to benzene, but in a highly excited state where it will mostly dissociate to CH + H.102 Recently, a new pathway via a smaller, five-membered ring radical (cyclopentadienyl Fig. 22) molecule has been proposed by Kaiser et al.354:
but this has yet to be added to photochemical models to assess its relative importance.
At higher pressures lower in the atmosphere, the three-body reaction combining two propargyl radicals becomes the dominant pathway for creation of benzene: 79, 80, 92, 102
Loss: Benzene is lost through ionization to the phenylium ion (CH) and through photolysis to form phenyl ()92, 357. The phenyl radical then either reforms benzene:94
or reacts with other radicals and neutral species, leading to molecules such as toluene (), styrene () and benzonitrile ().358
4.2 Nitrogen Compounds
Nitrogen compounds are formed by chemical combination of dissociation products from initial N and CH, and have formulas . All of the eight known heteroatomic nitrogen compounds are cyanides, wherein nitrogen is bonded to carbon by a triple bond (-CN) and therefore have a formula: -. These are HCN, HNC, CHCN, CHCN, CHCN, HCN, CHCN and CN. Other than the light molecules HCN and HNC, the remaining molecules are nitriles (organic cyanides). See Fig. 23.
Hydrogen isocyanide (HNC) is less stable than hydrogen cyanide (HCN) and is converted exothermically to HCN as it descends in the atmosphere. This leads to a predicted steep decrease in abundance with increasing pressure 359 and its present non-detection at lower altitudes.
In the lower atmosphere, nitrogen has always been found to date to be triple-bonded in the terminal position of a molecule: other types of species (amines, imines etc) have not yet been detected. We will return to the topic of what additional nitrogen compounds may be waiting to be discovered in a later section. The chemistry of known N-bearing molecules in the neutral atmosphere is now summarized.
4.2.1 Nitrogen
A major but unobserved constituent in Titan’s atmosphere was necessitated by the observed collisional broadening of methane spectral lines 360: this was hypothesized to be molecular nitrogen 361 which would be invisible at visible and longer wavelengths. The first conclusive observations of nitrogen were by Voyager 1’s UVS instrument, which detected dayside airglow at 96 and 98 nm, and longer wavelength absorptions with occultation measurements 362, 363. Measurements of nitrogen were greatly extended by Cassini’s UVIS instrument 26, 364, 365, 366, 367.
Production: The origin of nitrogen in Titan’s atmosphere has been long debated, and is not the subject of this paper. In brief, two major theories exist: enclathratization of N gas in the protosolar nebula368, or accretion in the form of NH ice followed by later photodissociation to eventually form N through a reaction cascade 369:
The latter scenario is currently favored due to the low temperatures in the sub-nebula required to capture molecular nitrogen directly. Variations on the theory include impact conversion of either NH or ammonium sulfate () to N 370, 371.
Loss: Molecular nitrogen is dissociated and/or ionized by short-wavelength solar radiation at nm 374, Saturn magnetosphere electrons 375 and Galactic Cosmic Rays (GCRs)177, 178, 179, 180 (Fig. 24)173.
and can recycle to through reaction with hydrogen, methane and other hydrocarbons, e.g. 242:
However, molecular nitrogen in the un-ionized state has very low reactivity, which in part contributes to its great abundance and significant longevity in the atmosphere.
Isotopes: Since and have significantly different UV cross-sections 377, it is important to correctly account for both isotopes and the wavelength variation of the solar spectrum to arrive at correct dissociation rates. Self-shielding by the more abundant is thought to reduce photolysis rates relative to the less abundant, less shielded , causing a lower ratio in nitrogen atoms than in the original molecules. Since significant amounts of atomic nitrogen go on to form nitriles, this skew towards increased production of may explain the lower in nitriles than in N itself 377, 378, 379.
Future directions: The dissociation and reaction pathways for N and its daughter ions and radicals remain one of the better known areas of Titan chemistry. However gaps remain, in particular whether nitrogen exists in chemicals such as amines and imines, or if it is incorporated in to heterocyclic ring molecules. This is further discussed in a later section.
4.2.2 Hydrogen Cyanide
Hydrogen cyanide was first detected by the Voyager 1 IRIS spectrometer through its strong infrared emission at 712 cm 380, and later at sub-millimeter wavelengths from ground-based observatories 381, 382, 383, 164. Although a relatively simple molecule that has been included in photochemical models for more than four decades, gaps in our knowledge of HCN formation may still exist, and new pathways have been identified recently 384.
Production: HCN is primarily produced in the upper atmosphere by the reaction of methane and nitrogen dissociation products (Fig. 25385):
and may be reformed from its ion by ion-molecule reactions, e.g.239:
HCN may also be lost in a two-step process, beginning with proton-transfer from a lower proton affinity molecule, e.g.:103
followed by dissociative recombination:373
Lower in the atmosphere, radical reactions and photolysis become important:91
As noted by previous authors, the CN triple bond is extremely stable and therefore the CN unit tends to persist when HCN is photolyzed, being incorporated into heavier nitriles, e.g.:
Also at low altitudes ( km 103) H-addition can lead to formation of the methylene-amidogen radical:
Future directions: Although well-studied for decades, recent work384, 388 has identified new pathways to the formation of HCN in planetary atmospheres for which reaction rates are currently unknown. Theoretical predictions now exist, but experimental confirmation is needed.
HCN has been shown to form co-crystals with hydrocarbons at Titan-relevant temperatures 389, the study of which will be important for understanding the solids and liquids on the surface. HCN, along with HCN, has also been implicated in the formation of CN in grain-surface chemical reactions 390, which requires further study to elucidate reaction rates and whether this process is sufficient to explain observed ice spectral properties 391.
Finally, HCN has been implicated in processes of astrobiological importance. A well known example is its proposed ability to directly form the amino acid adenine (CHN)392, 393, 394, 395 from the rearrangement (oligomerization) of five HCN molecules. Although the importance of this reaction for the seeding of life on the early Earth has been disputed396, 397, it may be more prevalent on Titan where HCN occurs in greater abundance.398, 399 HCN may also have the potential to polymerize into polyimines, structures that may catalyze astrobiologically important reactions 400. The astrobiological potential of HCN therefore remains under continued investigation401, 384, 402, 403.
4.2.3 Hydrogen Isocyanide
Hydrogen isocyanide, a higher energy isomer of hydrogen cyanide,404 was discovered on Titan using the Herschel space observatory by its sub-millimeter transition at 544 Ghz 405, and subsequently measured by ALMA as well 359, 128. HNC is readily interconverted to the more stable HCN (releasing kcal/mol),406 and therefore is predicted to have a steeply diminishing mixing ratio profile with altitude.98, 103
Production: HNC is produced by the same neutral reactions as HCN:
where the relative productions are estimated at 1300 km.90 At 1000 km R186 becomes dominant. Note that there are two important production pathways for :
with R188 dominating in the thermosphere and R189 becoming important in the mesosphere and below.90 Ion pathways may also be similar (see Fig. 26)385, although branching ratios are in most cases more uncertain than for HCN, e.g. through dissociative recombination of HCNH:407
HNC may also be produced as photodissociation product of CHCN387 in the upper atmosphere, and a further production peak may occur due to cosmic ray chemistry at 100-150 km 98.
Loss: At high altitudes (1300 km) the principal loss channels for HNC are:90
While at lower altitudes collisional isomerization to the lower energy HCN becomes important, and dominant by 600 km:90
Future directions: HNC/HCN is now one of two isomer pairs known in Titan’s atmosphere (the other being ). Study of the branching ratios and reaction rates leading to and from isomer pairs/triples etc is of importance because the less stable isomer(s) may follow different reaction pathways compared to the more abundant molecule(s). Therefore for a complete understanding of Titan’s atmospheric chemistry, all isomers must be included in models. Study of the vertical ratio between HCN/HNC and CHCCH/CHCCH may also provide useful information on the abundance of atomic H, collisions which can cause conversion between the isomers.
4.2.4 Acetonitrile
Acetonitrile was first detected on Titan in the early 1990s by millimeter wavelength astronomy 408, followed ten years later by the first measurement of its vertical profile 164 using the 30 m telescope at IRAM. CHCN was the first Titan molecule to be first detected at millimeter wavelengths, an astronomical technique that was to yield many other discoveries later with ALMA.
Production: Acetonitrile is produced in the upper atmosphere by the reaction of N-radicals with ethylene409:
and by the termolecular reaction of H with cyanomethyl (CHCN), in a chain that begins with acrylonitrile (CHCN):98, 103
Loss: The major loss mechanism for acetonitrile is proton transfer from another ion to form :410
Future directions: Acetonitrile, as many other simple molecules, has been implicated in formation of a co-crystal with acetylene 417, providing an interesting avenue for further investigation of its solid phase properties, with possible implications for cloud particle growth.
4.2.5 Cyanoacetylene
[…]
Production: Cyanoacetylene is produced above 1000 km by reaction of the CN radical from photolysis of HCN with acetylene (see Fig. 28)202, 98, 419, 203:
and to a lesser extent by photodissociation of acrylonitrile (see Fig. 30).
Loss: As with other nitriles, the principal loss pathway for cyanoacetylene in the upper atmosphere is proton transfer, forming , e.g.:239
followed by dissociative recombination breaking up the molecule 412:
On the other hand, photolysis is not a significant loss channel, since is thought to rapidly recycle back to through reaction with methane:98
While HCN does react without a barrier with radicals such as CN and CH, the main loss channel in the neutral atmosphere is thought to be successive hydrogen addition:98
Future directions: In interstellar space ( molecular clouds such as TMC-1), cyanopolyynes of the form have been detected with 420, 421, 422. has been sought, but not yet detected, in Titan’s neutral atmosphere. Detection of this molecule may provide some clues as to the relative abundances of cyanopolyynes vs N-heterocycles.
4.2.6 Cyanogen
Cyanogen (CN), like cyanoacetylene was first detected in Titan’s atmosphere by the Voyager IRIS spectrometer in the infrared 342 at 233 cm.
Production: Cyanogen is thought to be produced mainly by addition of CN to HNC:
and through the radical-radical reaction:98
via the intermediate adduct NCHCN. Neither of these reactions are expected to have an entrance barrier98, 423, 424, while the reaction of CN with HCN is inefficient due to the low rate constant.425, 426, 427
Loss: Cyanogen is lost by photodissociation (Fig. 29428) and by H-addition (with an entrance barrier of 14-30 kJ/mol):98
[…]
4.2.7 Acrylonitrile
Acrylonitrile (CHCN) was the second molecule to be discovered on Titan at millimeter wavelengths using ALMA 166, following the detection of propionitrile,167 discussed in the next section.
Production: Acrylonitrile (see Fig. 3098, 415, 431, 432) is produced above 800 km by substitution of the CN radical onto ethylene202, 203:
Loss: In a similar manner to HCN, and other nitriles, CHCN is lost in the ionosphere by the two-step process of proton transfer:433, 434
followed by dissociative electron recombination:435
In the lower atmosphere it may be lost to photodissociation (which tends to recycle acrylonitrile) or by H-addition:94
Future directions: Several small nitrile molecules, which tend to exhibit polar properties, have been investigated in a theoretical study for the potential to self-organize into spherical vesicles or membranes in non-polar liquids (such as, for example, Titan lakes and seas of methane-ethane-nitrogen). These calculations showed that acrylonitrile was the best candidate for forming so-called ‘azotosomes’ 436, which, if experimentally confirmed, could be significant for astrobiology, as vesicles (containers) for self-replicating organisms. However at this time experimental verification of azotosomes is still lacking, while a later study has questioned the ability of these structures to form 437.
4.2.8 Propionitrile
Propionitrile (CHCN) was the second molecule to be originally detected using sub-millimeter astronomy and the first molecule with ALMA 167.
Production: Propionitrile has been posited to be produced above 900 km 103 by the association reactions (see Fig. 31)98, 438:
In the middle atmosphere (400–900 km) 103, successive rounds of hydrogen addition to acrylonitrile via termolecular reactions can produce propionitrile:
or the termolecular reaction:98
Loss: As with other nitriles, the first step in loss of this nitrile in the ionosphere is proton transfer, forming :
This is followed by either dissociative electron recombination:440
or ion-neutral reactions such as:323
Future directions: Propionitrile has been asserted to condense in pure crystalline form in Titan’s atmosphere 441, on the basis of an unexplained feature in Titan’s far-infrared spectrum. This has been questioned based on vapor pressure of the gaseous form 442, 443, although it is possible that a co-condensed ice containing CHCN along with other gases may replicate the unexplained ‘haystack’ emission 444. Further work on spectroscopy will be required to determine if this is a unique solution, or if other possibilities exist.
4.2.9 Cyanopropyne
Cyanopropyne CHCN was the fourth molecule to be discovered by ALMA spectroscopy of Titan at mm wavelengths 168, following previous detection in the ISM 445.
Production: Production pathways for cyanopropyne (see Fig. 32) are more uncertain than for many other molecules due to the size and complexity of the molecule, allowing for more numerous reaction pathways, and multiple isomers of . Pathways involving radicals include CN substitution onto propyne or butadiene 446, 447:
or attack on ethylene:448
Loss: Cyanopropyne is thought to be lost through either photolysis, or through protonation, e.g.:103
followed by dissociative electron recombination:103
Future directions: , has at least three stable isomers that have been detected in space 450, 451. Besides the currently detected isomer (cyanopropyne, CHCN, butynenitrile or methylcyanoacetylene), there is also cyanoallene ()452 and propargyl cyanide ()453, both being first detected in the Taurus Molecular Cloud (TMC-1) at radio wavelengths. These provide good targets for detection on Titan, and their measurement would help to constrain photochemical pathways and models. Further more exotic arrangements of the same atoms may also exist and remain to be detected.
4.3 Oxygen Compounds
[…]
4.3.1 Water
Water was first detected in Titan’s atmosphere through infrared spectroscopy with ISO 457 through detection of emission lines at 39.4 and 43.9 m, and subsequently confirmed with Cassini CIRS 458, 459.
Production: Water (Fig. 34460, 461, 462) is thought to mainly be derived by the re-combination of OH infalling at the top of the atmosphere, primarily sourced from dissociated Enceladus water, with methane and its dissociation products :
Loss: Water is lost to photolysis throughout the atmosphere, reforming hydroxyl (OH). A large fraction of this OH reacts with to reform water (see previous equation). However, OH participates in several other reactions. Above 1000 km, it reacts with to form NO:97, 373
while in the middle atmosphere it reacts with CO to form CO: 97
and any remaining unreacted water is ultimately lost by condensation in the lower stratosphere.
Future directions: Due to its low vapor pressure, water remains difficult to measure in Titan’s atmosphere. Currently there are large uncertainties in its vertical profile 458, 459, and its latitudinal distribution remains unknown. Further work to better constrain these distributions may help to elucidate the relative importance of meteoritic versus Enceladus sources 97.
4.3.2 Carbon Monoxide
Carbon monoxide (CO, Fig. 35464) was first detected on Titan by near-IR spectroscopy, showing an absorbance of CO at 1.6 m 465, and the detection was soon confirmed at radio wavelengths 381. Estimates of its abundance fluctuated throughout the years following its discovery 466, and because these measurements were often sensitive to different altitudes, this led to the suggestion that the vertical profile was non-uniform 381, 467. Subsequent measurements with high sensitivity telescope arrays at Owens Valley and Mauna Kea however showed evidence for a uniform profile, converging on a mixing ratio of ppb 468, 469.
Recently, high-sensitivity observations with ALMA have narrowed the experimental error to a range of ppm 133, making it the fourth most abundance species in Titan’s atmosphere after N, CH and H. As a triple-bonded molecule, CO once produced is both resistant to photolysis and chemical reaction, as well having no loss through condensation, and hence all evidence available at present points to a uniform vertical profile.
[…]
(Formaldehyde may also be created by reaction of OH with CH, CH etc).
(also OH + CH, OH + CH etc). Note that formaldehyde has yet to be detected in Titan’s atmosphere.
Lesser routes to CO production may be through reaction of atomic oxygen (deposited to the top of the atmosphere from Enceladus) with methane fragments, e.g.:81
Loss: CO is primarily lost slowly through reaction with OH forming CO:474
which is in turn lost through condensation.
Future directions: Laboratory chemistry simulations of Titan’s atmosphere have shown that CO may react with methane and nitrogen when sufficiently stimulated, forming amino acids and even nucleobases 475. This provides an exciting possibility of astrobiology that now requires remote and in situ measurements to confirm.
4.3.3 Carbon Dioxide
Carbon dioxide (CO) was one of seven new gas species detected by Voyager’s IRIS infrared spectrometer 476, 477 and subsequently by Cassini CIRS 478, 278. Unlike shorter-lived chemical chemical species ( HCN, CH), CO exhibits little variation with latitude in the lower stratosphere, lacking a polar enhancement.
Production: CO is thought to be mostly produced from CO + OH as shown in the previous section (R 255). See also Fig. 36464, 479, 103.
and also through condensation in the lower stratosphere.
4.4 Growth of large particles
As hydrocarbon molecules grow to ever-larger sizes, they may take several forms: long chains, fused rings (PAHs - polycyclic aromatic hydrocarbons), or rings connected by hydrogen bonds (poly-phenyls) (see Fig. 37). Nitrogen incorporation is also likely, for example in the form of PANHs (polycyclic aromatic nitrogen heterocycles). It is thought that eventually larger molecules clump together due to electrostatic forces to form fractal aggegrates 486, 212, 487, 488 - Titan haze particles which form the well-known golden haze at visible wavelengths. These in turn become the nuclei for stratospheric hydrocarbon ice particles or tropospheric methane raindrops 114, 115, 116, 117, and fall to the surface where they form Titan’s dune fields 118.
[…]
Production: The formation of PAHs by addition to benzene rings has been a topic of debate for considerable time. Recently, Kaiser and Hansen353 have categorized possible pathways into five principal mechanisms:
HACA - "hydrogen abstraction- CH addition"490
HAVA - "hydrogen abstraction- vinylacetylene addition"
PAC - "phenyl addition-dehydrocyclization"
RRR - "radical-radical reactions"
MACA - "methylidene addition-cyclization aromatization"
Each of these offers potential pathways to larger molecules. In brief, the HAVA mechanism, in which vinylacetylene () adds to aromatic rings (such as benzene) in a barrierless reaction is thought to be principle mechanism by which additional six-membered rings are added to existing rings at low temperatures, such as in planetary atmospheres. The other mechanisms offer alternate routes to addition of five and six-sided rings, predominantly at high temperatures of 1000s of K (HAVA, PAC, RRR), although MACA may operate at low temperatures to form indene. The reader is directed to the review paper by Kaiser and Hansen353 for a full description, which is beyond the scope of this article.
No discussion of aerosol particle growth would be complete without mention of negative ions. The discovery of large negatively charged ions at high altitudes by Cassini’s CAPS instrument was one of the major surprises about Titan’s atmosphere early in the mission 61, 60, 69, 491, 492, 68. Small negative ions may be formed by several processes, including dissociative electron attachment, e.g.:
and by radiative attachment to a radical species already formed through photochemistry:103
Once H is produced, it leads to the creation of some larger negative ions through proton abstraction:103
Successively larger aerosol particles are produced through a variety of ion-neutral and ion-ion reactions 493, 494, 495. The largest charged particles tend to be predominantly negative ions, due to the larger to their higher mobility.
Loss: Since aromatic rings, once formed, are very stable, with many possibilities to disperse absorbed energy internally, their principle loss channels will be either (a) to form radicals and then larger molecules; (b) to agglomerate or (c) to condense.
Future directions: Much work is still required to further elucidate the growth of large particles, especially the relative important of ion vs neutral chemistry at different altitudes. See also the later Section on the topic of PAHs.
5 Gaps in our knowledge
In this section we consider where the gaps are in our current knowledge and understanding of Titan’s atmospheric composition and chemistry, and how these gaps might be addressed in the near future through combinations of astronomy, laboratory and theoretical work.
5.1 Aliphatic Species
5.1.1 Hydrocarbons
At present, CH remains the only hydrocarbon species definitively detected in the neutral atmosphere, while no aliphatic species with have been detected (benzene, c-CH, being detected as a ring).
Species with four or more carbon atoms show increased possibilities for structural isomerization, as illustrated in Fig. 39. This results in both increased challenges for observational detection, as well as new and interesting chemical possibilities such as branched chain molecules.
Upper limits for -butane and -butane (CH) have been calculated as and respectively, from CIRS spectra at 200–250 km, 30N–30S498.
Future Work: Astronomical detection of species such as , (i.e. butynes and butenes) and even would greatly help to improve constraint on photochemical models, which are currently lacking in data to model for this regime. Building our knowledge of aliphatic species such as C4’s and C5’s may provide a better understanding of the pathways to benzene or other cyclic species.
In parallel, photochemical models that currently treat multiple isomers under single formulae such as and must continue to expand treatment of separate isomers. To date this has been sparse and primarily for a few specific cases for which isomeric data exists: especially HCN and HNC, and CHCCH and CHCCH, which are considered separately in current models 90, 91, 99, 103. A good reason for lack of inclusion of separate isomers is lack of knowledge of branching ratios and isomer-specific reaction rates. Monte Carlo simulations have proved useful at showing which reaction rate uncertainties have the biggest effects on the uncertainties in the solution 87, 88. These researches thereby provide important prioritization of where resources such as laboratory time and theoretical effort (e.g. TST calculations) can best be spent to most rapidly improve our knowledge.
5.1.2 Nitrogen Species
All of the nine detected nitrogen compounds in Titan’s neutral atmosphere, formed from dissociation products of N and CH, are triple bonded. These include N itself, and eight known cyanides, which include nitrogen in a terminal -CN functional group.
[…]
| crotonitrile | () | |
| butanenitrile | () | |
| isobutyronitrile | () | |
| cyanocycloproane | () | . |
Nitrogen has yet to be detected in other types of bonding - i.e. where it is not in a triple bond, such as in amines, imines (beyond HNC), azines and nitrogen heterocycles (see Fig. 40). A simple example is ammonia (NH) which was tentatively inferred from Cassini’s mass spectra at high altitudes, at mixing fractions of , although not yet uniquely detected due to the barometric degeneracy problem of unit resolution mass spectroscopy.501, 219
The major channel for ammonia production is thought to be via NH + CHN, as follows:103
Ammonia is mostly lost through photodissociation.
Upper limits for some molecules in the lower atmosphere have been estimated, including for ammonia (NH) of 1.3 ppb (3-) at 107 km, 25S 502, and methanimine (CHNH) at 0.35 ppb (3-) in the stratosphere 503.
Future Work: No molecules having both oxygen and nitrogen, e.g. HNO, and more complex amino acids, have been detected. Detection of such functional groups and molecules, important for biological activity on Earth, is a key area of future research. Several nitriles have been detected only through sub-millimeter astronomical techniques as previously mentioned, but not using infrared techniques despite several decades of attempts 499, 500, 49, 502. In fact laboratory line lists with intensities do not exist for many of the nitriles sought in the infrared, except CHCN 504, 505, hampering the search and implying a need for new laboratory work to obtain cold temperature spectra.
5.1.3 Oxygen species
Few oxygen compounds have been detected in Titan’s atmosphere: only CO, CO and HO to date. Diatomic molecular oxygen, , readily found in the atmospheres of the inner planets, is absent, allowing organic chemistry to proceed to great complexity. The oxygen compounds detected appear to be attributable to an external source of oxygen (both O and OH), most likely originating from Enceladus 57, 81. Prior to the discover of the Enceladus plumes a meteoritic source was favored 506, and some meteoritic contribution may still be present 97.
[…]
Photochemical models 97 predict that trace amounts of molecules such as HNO, HNCO, HCO and CHOH should be present in the atmosphere, and perhaps detectable at high altitudes by observatories such as ALMA. To date, few published studies have attempted to directly identify further oxygen species. The Cassini INMS team published upper limits for methanol () and acetaldehyde () of 30 ppb and 10 ppb in the ionosphere at 1100 km 219. In the stratosphere, upper limits have been determined for methanol and formaldehyde of 6 ppb and 2 ppb respectively at at 107 km, 25S 502 from infrared spectroscopy with Cassini CIRS.
Future Work: Vibrational (IR) and/or rotational (sub-mm) line lists exist for most small oxygen compounds currently undetected in Titan’s atmosphere (see Fig. 41), making astronomical searches viable.509, 510, 511
5.2 Cyclic molecules
5.2.1 Single hydrocarbon rings
Two small cyclic molecules have been definitively detected in Titan’s neutral atmosphere: cyclopropenylidene (c-CH) 165 and benzene (c-CH) 49. Other small cyclic molecules likely to exist include the saturated cycloalkanes - cyclopropane (), cyclobutane () and larger - and possibly cycloalkenes such as cyclopropene (), cyclobutene (), cyclobutadiene (), and others. Substituted rings are also possible: cyanocyclopropane ().
In the ISM, several single ring molecules have been detected. These include the small 3-carbon rings cyclopropenylidene (c-CH 306) and its related radical 512, and a substituted species, ethynyl cyclopropenylidene () 513. The five-sided ring cyclopentadiene (), and the six-sided rings benzyne () and benzene (c-CH) have also been detected 348, 513, 514.
Future work: Work is needed on all fronts to advance our understanding of the formation pathways, stability and prevalence of small cyclic rings in Titan’s atmosphere (see Fig. 42). These include astronomical observations, laboratory work and photochemical modeling. In the 2030s, we anticipate direct in situ measurement of such molecules by the Dragonfly probe DraMS (Dragonfly Mass Spectrometer) instrument, which, unlike Cassini INMS will have ability to definitively identify molecular structure through a combination of tandem mass spectrometry (MS/MS) and Gas Chromatograph Mass Spectrometry (GCMS) 515.
5.2.2 Multi-ring hydrocarbons: PAHs and polyphenyls
Polycyclic aromatic hydrocarbons (PAHs) are multi-ring molecules composed of carbon and hydrogen that exhibit aromatic character, that is to say they have delocalized electron bonds. Example include naphthalene (two 6-membered rings), indene (one five-sided and one six-sided ring), anthracene and phenanthrene (three 6-membered rings), and larger examples (see Fig. 43). Benzene is not considered to be a PAH, since it has only one ring.
PAHs have long been suspected to exist in interstellar space516, and have been implicated as culprits responsible for the so-called ‘diffuse interstellar bands’ (DIBs)517 (see review by 518). To date, only a single non-functionalized PAH has been uniquely identified in space (indene - 513), although a greater number of CN-substituted single and double rings (cyano-PAHs) have been identified, assisted by their strong rotational lines due to the cyanide group - see review by McCarthy and McGuire519.
[…]
Laboratory experimental work on haze formation by UV photolysis has shown that the inclusion of benzene in initial reagent mixtures along with N and CH leads to the formation of significantly larger molecules than when a N/CH mixture is used 523. More recently, naphthalene has also be used as a starting reagent in lab tholin experiments 524 showing similar results. It should be noted that, on Titan, both benzene and naphthalene would presumably first have to form from methane, so their use in lab experiments may be considered an acceleration of a natural process that could in principle start from a pure N/CH mixture and arrive at the same result over long time periods.
Multi-ring organic molecules are not constrained to form as ‘fused’ ring PAHs such as naphthalene, but may instead form as polyphenyls (see Fig. 37). It has been argued that a significant amount of carbon rings in Titan’s atmosphere may be in the form of polyphenyls rather than as fused rings 525. Titan’s aerosols are likely to be a mixture of fused and unfused rings, forming monomers and then fractal aggregates 526.
Future work: Future laboratory, and eventual in situ experimental work is required to determine the relative importance of fused vs non-fused rings.
5.2.3 Fullerenes
Fullerenes are carbon allotropes formed of rings of 5 to 7 atoms in closed or partially closed mesh structures. These can include ‘buckyballs’ (or buckminsterfullerenes), such as the spherical () and ellipsoidal () molecules 527, 528, but also as sheets (graphene) and cylinders (carbon nanotubes) (see Fig. 44). Despite their large size and apparent complexity, buckyballs have been detected in space both as neutrals 529 and ions 530, and in meteorites 531. Fullerenes have been hypothesized to exist in Titan’s atmosphere 532 although a recent attempt to detect them in Spitzer data proved unsuccessful.533
Future work: Greater sensitivity with observatories such as JWST 534 may enable more sensitive searches for fullerenes in Titan’s atmosphere. In addition, little lab work has been done at present to determine what effect the presence of fullerenes could have on Titan’s atmospheric chemistry, aerosol formation and surface geology.
5.2.4 PANHs
Nitrogen heterocycles and PANHs (polycyclic aromatic nitrogen heterocycles) are similar to PAHs, but with nitrogen incorporation into the ring structure (see Fig. 45, 165). Nitrogen-heterocycles have been sought unsuccessfully in interstellar space, with upper limits for molecules such as pyridine and quinoline derived 535. Mass spectroscopy of Titan’s atmosphere with Cassini INMS has identified peaks at masses that could correspond to N-heterocycles, such as (could be protonated pyridine) at mass 80 and at mass 81 (possibly protonated pyrimidine) 219. However, aliphatic variants are possible making the PANH ion identification uncertain.
[…]
Laboratory work has examined how tholin (Titan haze analog) formation in UV photolysis experiments is affected by the inclusion of N-heterocycles such as pyridine and quinoline 536, 537. Changes in the spectrum show similarities to features in Titan’s haze spectrum 536, and the structures formed show a mixture of polymeric and random co-polymeric structure 524.
Future work: In future, more sensitive astronomical searches may be undertaken, for example using ALMA and JWST. Photochemical models at present do not include detailed description of N-heterocycle formation and growth, which must be included in future generations to adequately model processes leading to haze formation.
5.3 Sulfur and Phosphorous Chemistry
[…]
In principle the O/S ratio should allow further constraint of the source of Titan’s oxygen flux, since the O/S ratio is predicted to be some 1000 less for a cometary source (O/S100)542 than an Enceladus source (O/S)543. Detection of sulfur in Titan’s atmosphere may also provide evidence for cryovolcanic activity 252.
Both phosphorus and sulfur are among the six most essential elements for biochemistry on Earth (the so-called CHNOPS elements). With four of the six already detected on Titan, it is therefore of considerable interest to seek the remaining two, to further assess Titan’s potential for astrobiology. Recent reports that all six CHNOPS elements have now been detected in the plume material of Enceladus 543, 544, make it feasible that some trace amounts of P and S arrive at the top of Titan’s atmosphere, as is the case apparently with O57.
Future work: Future work is required on both the direct detection of P and S-containing substances by both astronomical and in-situ techniques - but also in laboratory work, computer photochemical modeling, and clarification of reaction pathways and rates, especially at low temperatures.
5.4 Radicals
For completeness, we include consideration of radical species - atoms and molecules with unpaired free electrons - even though these are highly reactive and unlikely to be found in significant numbers, or significantly deep in Titan’s neutral atmosphere. Radicals include fragments of CH, such as CH (methylidene), CH (methylene), and CH (methyl), as well as ground state and excited nitrogen atoms formed from break up of N, such as as N(D) and N(S).
Many radicals (CN, OH etc) have been observed in the ISM 450, including at least one cyclic radical ()512; also in comets 545 and tenuous satellite exospheres 546. However, to date only the methyl radical has been detected by astronomical techniques in a planetary atmosphere 547, 548.
Future Work: ALMA observations are particularly sensitive to Titan’s upper atmosphere, and can sense HCN at altitudes of up to 1200 km 128, 159. Future investigations with ALMA may allow the detection of radicals with dipoles, and more sensitive infrared observatories such as JWST may prove effective at detecting radicals such as methyl in the infrared 534.
6 Conclusions
The organic-rich atmosphere of Titan constitutes the most complex atmospheric chemical network known outside of Earth. This provides a unique natural laboratory for understanding the synthesis of organic compounds, processes that may have been important early in the history of the Solar System 549, 550, and may have seeded the origins of life on Earth 551, 552.
Therefore it is of substantial scientific importance to better understand these processes and chemical results. This field of enquiry brings together astronomers, laboratory chemists, theoretical chemists and atmospheric modelers whose combined approaches are needed to unravel the entire picture. Substantial progress has been made, especially stimulated by the recent wealth of data from the Cassini-Huygens mission 2, 3, and the selection of the Dragonfly mission275 that will arrive in the 2030s.
Astronomy: Titan is a distant object and gathering robust information is difficult in the absence of spacecraft: astronomy is currently the only means to gather data about Titan directly. Currently active ground and space-based observatories such ALMA, IRTF and JWST are providing continuity of data collection since the end of Cassini-Huygens regarding seasonal changes in Titan’s atmosphere,159, 147 and new information on the chemistry, composition and isotopic ratios.170, 171, 325, 553, 166, 168, 165, 554, 133.
Laboratory Studies: Whilst data collection through astronomy and remote sensing continues, a robust campaign of laboratory experiments and theoretical work is continuing in parallel to understand the origins and chemical evolution of the atmosphere, and interaction with the surface and subsurface. These diverse enquiries include spectroscopy of gases,555, 341, 556, 557, 558, 559, 560, 498, 561, 562, 563 measurement of reaction rates,358, 347, 564, 565, 409, 242, 566, 567, 354, 568 experimental work on co-crystals,297, 276, 277, 417, 569, 570 ices,571, 572, 573, 574, 575, 576, 577, 578, 579, 580, 581, 582 and hazes (tholins).111, 110, 583, 584, 585, 586, 587, 588, 589, 590, 591, 592
Modeling: The chemistry of Titan’s atmosphere cannot proceed without disturbing the medium in which it takes place: the minor gases and haze generated have a substantial effect on the heating and cooling of the atmosphere 593, and in turn changes to the thermal structure of the atmosphere leads to dynamical motions including vertical eddy mixing, and meridional transport. As gases are transported, they enter atmospheric regions that may have greater or lesser photon flux from the Sun, have fewer or greater opportunities to interact with electrons, and encounter differing densities of radicals and other reagents. Therefore, decoupling chemistry from dynamics is not possible. Combining the current generation of Global Circulation Models (GCMs) 594, 595, 155, 596, 597, 598 and photochemical models 99, 100, 102, 103, 599 to create 2D and 3D coupled chemical-GCMs is a challenging but important task that must occupy the next generation of modelers.
Future Missions: The Dragonfly mission estimated to land on Titan 2034 will provide a wealth of new data about Titan’s surface and atmospheric boundary layer. However Dragonfly will only investigate the low latitudes, including dunes and the crater Selk. In the future an orbiter, as envisaged by several published studies 600, 601, 602, would provide the important benefit of a truly global picture, including potentially complete global mapping of the atmosphere and surface at uniform resolution, with time-domain information to search for changes occurring. Other elements such as a balloon, airplane and/or floating lake probe could provide valuable in situ information about other environs 603, 604, 605.
There is no doubt that Titan still offers many challenges to our understanding that will provide fertile areas of study for future generations of scientists 21, 606, and offer rewards through important insights into the chemical evolution of the Solar System and origins of life in the universe.
Funding for this work was through NASA’s Astrobiology program. The author wishes to thank the Astrobiology Program manager Dr Mary Voytek (NASA HQ) and the Principal Investigator of the CAN-8 Project "Habitability of Hydrocarbon Worlds: Titan and Beyond" Dr Rosaly Lopes (JPL/Caltech) for their support of this work. Thanks is also due to Nicholas Lombardo and Alexander Thelen for providing text files of retrieved gas profiles from prior publications. The author sincerely thanks two anonymous reviewers and the guest editors of the ACS Earth and Space Chemistry special edition on astrochemistry, Martin Cordiner and Christopher Bennett for their thoughtful comments and feedback that helped to improve the manuscript; and to the Journal Editor Eric Herbst, and staff for their assistance in the review and publication process. Last but not least, the author is very grateful to a cadre of early-career students and postdocs who proof-read parts of the final submitted version of the manuscript: Brandon Coy (UCLA), Nicholas Kutsop (Cornell University), Paige Leeseberg (University of Iowa/SURA), Siobhan Light (University of Maryland/SURA), Nicholas Lombardo (Yale University), Edward Molter (University of California Berkeley), Jonathon Nosowitz (Catholic University), Alexander Thelen (Caltech). Any remaining errors or inaccuracies are the responsibility of the author.
Appendix A Acronyms and Abbreviations
| ALMA | Atacama Large Millimeter/submillimeter Array |
| CAPS | Cassini Plasma Spectrometer |
| CDA | Cosmic Dust Analyzer (Cassini instrument) |
| CIRS | Composite Infrared Spectrometer (Cassini instrument) |
| DR | Dissociative Recombination |
| DSMC | Direct Simulation Monta Carlo |
| ESA | European Space Agency |
| GCM | Global Circulation Model |
| GCMS | Gas Chromatograph / Mass Spectrometer (Huygens instrument) |
| HASI | Huygens Atmospheric Structure Instrument |
| INMS | Ion and Neutral Mass Spectrometer (Cassini instrument) |
| IRAM | Institut de Radioastronomie Millimetrique |
| IRIS | Infrared Interferometer Spectrometer (Voyager instrument) |
| IRTF | Infrared Telescope Facility |
| ISM | InterStellar Medium |
| ISO | Infrared Space Observatory |
| IUPAC | International Union of Pure and Applied Chemists |
| JWST | James Webb Space Telescope |
| LTE | Local Thermodynamic Equilibrium |
| NASA | National Aeronautics and Space Administration |
| PAH | Poly-Aromatic Hydrocarbon |
| PANH | Poly-Aromatic Nitrogen Heterocycle |
| RPWS | Radio and Plasma Wave Spectrometer (Cassini instrument) |
| RSS | Radio Science Subsystem (Cassini| instrument) |
| SNR | Signal to Noise Ratio |
| TEXES | Texas Echelon Cross Echelle Spectrograph |
| TMC | Taurus Molecular Cloud |
| TST | Transition State Theory |
| UVIS | Ultraviolet Imaging Spectrometer (Cassini instrument) |
| VIMS | Visible and Infrared Mapping Spectrometer (Cassini instrument) |
| VMR | Volume Mixing Ratio |